### v1

# Robot Navigation with Entity-Based Collision Avoidance using Deep Reinforcement Learning

Abstract
Efficient navigation in dynamic environments is crucial for autonomous robots interacting with various environmental entities, including both moving agents and static obstacles. In this study, we present a novel methodology that enhances the robot’s interaction with different types of agents and obstacles based on specific safety requirements. This approach uses information about the entity types, improving collision avoidance and ensuring safer navigation. We introduce a new reward function that penalizes the robot for collisions with different entities such as adults, bicyclists, children, and static obstacles, and additionally encourages the robot’s proximity to the goal. It also penalizes the robot for being close to entities, and the safe distance also depends on the entity type. Additionally, we propose an optimized algorithm for training and testing, which significantly accelerates train, validation, and test steps and enables training in complex environments. Comprehensive experiments conducted using simulation demonstrate that our approach consistently outperforms conventional navigation and collision avoidance methods, including state-of-the-art techniques. To sum up, this work contributes to enhancing the safety and efficiency of navigation systems for autonomous robots in dynamic, crowded environments.
Keywords Robotics  $\cdot$ Reinforcement learning  $\cdot$ Collision avoidance  $\cdot$ Deep learning  $\cdot$ Robot navigation  $\cdot$ Motion planning
1 Introduction
The development of robots, designed to assist humans and perform different tasks, has revolutionized numerous sectors, including industrial automation, manufacturing, agriculture, space, medical science, household utilities, delivery, and social services. Autonomous navigation, a key attribute of these robots, enables them to move independently, perceive the environment, and take appropriate actions to complete tasks efficiently, effectively, and safely.
Recent research interest in mobile robots, particularly in the context of practical applications such as delivery, search and rescue, service, and warehouse robots has emphasized the importance of robot navigation. This field solves the problem of planning the robot’s path from the start point to the destination while avoiding other agents and obstacles in different dynamic environments. This task becomes even more challenging when the robot needs to navigate through crowded places like airports, hospitals, streets, squares, and shopping centers. Traditional navigation approaches, which often treat moving agents as static obstacles or react through short-sighted, one-step look-ahead, result in unsafe and unnatural behaviors.
The necessity for a robot to navigate in a socially compliant manner and understand the dynamic nature of its environment, especially in populated areas, is a challenging task. Furthermore, it is crucial for the robot to consider the diversity of environmental entities around it, as each entity exhibits unique characteristics, patterns of movement, safety requirements, and consequences of a collision. In this work we use the term “environmental entities", or simply entities, to collectively describe both the moving agents and the static obstacles presented in the environment. To get examples of moving agents let us consider sidewalk autonomous delivery robots. These robots are primarily used for last-mile deliveries and typically operate on pavements, courtyards, pedestrian crossings, and occasionally on parts of the road. In this case, robots interact with such moving agents as adults, children, bicyclists, scooters, vehicles, motorcycles, etc. Static obstacles in this case are curbs, trash cans, trees, bus stops, buildings, etc. For the sake of brevity, in what follows we will sometimes omit the word static, and by “obstacles" we will mean static obstacles.
For robots to navigate in densely populated spaces in a socially compliant manner, it is crucial to understand the behavior of surrounding agents and obstacles and adhere to cooperative rules. To achieve that, we implement a new method, based on deep reinforcement learning, which allows the robot to navigate in dynamic environments and to interact differently with different types of agents and obstacles based on various safety requirements. We call this approach Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL).
Our main contributions can be summarized as follows:
  • We propose a new method of collision avoidance based on deep RL which can effectively utilize information about the types of entities, which helps the robot to avoid collisions and leads to safer interaction and navigation.
  • We propose a reward function that penalizes the robot differently depending on the type of agent or obstacle that the robot collides with. In our research, we consider such types of agents and obstacles: adults, bicyclists, children, and static obstacles. Apart from that, our new reward function penalizes the robot for being close to entities and the safe distance also depends on the entity. Moreover, our reward function encourages the robot not only to achieve the goal but also to be close to the goal.
  • We provide an improved algorithm for training and testing the model which significantly improved the speed of training and testing and allowed us to train the robot in complex environments.
  • We test the applicability and the effectiveness of the proposed approach in our simulator. Our experiments demonstrate that the new approach consistently outperforms standard methods of robot navigation with collision avoidance including the state-of-the-art methods.
The rest of the paper is organized as follows: in Section 2.1 we give an overview of the related work; in Section 2.2 we formulate the problem; in Section 3 the proposed EB-CADRL approach is presented, including the description of the reward function, model and training algorithm; Section 4 is dedicated to the experimental part, evaluation methodology, in this section we provide the results and ablation experiments; in Section 5 we discuss the limitations of our work and provide ideas for future work; the conclusions are made in Section 6.
2 Background
2.1 Related Work
Recently extensive research has led to significant advancements in robot navigation including navigation in crowded environments.
Social Force Model (SFM) Helbing and Molnár (1995) uses attractive and repulsive forces to model crowd interactions. Other approaches like Reciprocal Velocity Obstacle (RVO) and Optimal Reciprocal Collision Avoidance (ORCA) van den Berg et al. (2011) consider surrounding agents as velocity obstacles to calculate optimal collision-free velocities under reciprocal assumptions. These methods, while effective in specific scenarios, rely heavily on hand-crafted rules and often fail in real social scenarios where their assumptions are not fully met, compromising the safety of human-robot interaction and leading to the “freezing robot" problem in dense environments.
Recently, learning-based methods have gained significant attention. Several studies use imitation learning to derive navigation strategies from demonstrations of desired behaviors. In studies Tai et al. (2018), Long et al. (2016) navigation policies using depth images, lidar measurements, and local maps as inputs are developed by imitating expert demonstrations. In addition to behavioral cloning, inverse reinforcement learning has been applied in Kretzschmar et al. (2016) to extract the underlying cooperation features from human data using the maximum entropy method. However, the effectiveness of these methods heavily relies on the scale and quality of the demonstrations, which can be resource-intensive and limit the policy’s quality due to human constraints.
Another approach treats robot social navigation as a Markov decision process and explores it through deep reinforcement learning, considering human safety and comfort Chen et al. (2017a). Chen et al. Chen et al. (2017b) introduce a decentralized multi-agent collision avoidance algorithm that employs a value network encoding the estimated time to the goal based on the joint state of the robot and surrounding agents. Following this, Everett et al. Everett et al. (2018) propose an enhanced algorithm using Long Short-Term Memory (LSTM) to sequentially process the states of neighboring agents, enabling navigation among an arbitrary number of agents. However, the network models in these methods are relatively simplistic and struggle to account for the crowd’s potential impact, hindering safe human-robot interaction. In contrast, Chen et al. Chen et al. (2019) present a value network that simultaneously models human-robot and human-human interactions to enhance the robot’s decision-making. Nonetheless, this method captures the spatial relationships of pedestrians but overlooks the temporal relationships of their movements, limiting the robot’s ability to predict pedestrian trajectories. In Xue et al. (2024) authors propose a deep reinforcement learning framework with a value network for mobile robot navigation that leverages spatial-temporal reasoning to understand crowd interactions, designs hazardous areas based on pedestrian speeds, and formulates a reward function to ensure pedestrian safety and comfort.
In our work, we aim to address the limitations of the methods mentioned above by creating a more robust, safe, and socially compliant navigation method using deep reinforcement learning.
2.2 Problem Formulation
We consider the task of a delivery robot navigating towards a goal in a dynamic environment populated with adults, children, bicycles, and static obstacles. This task can be formulated as a partially observable Markov decision process (POMDP). The robot must navigate in a socially compliant manner, avoiding collisions while reaching its goal. We define the state of each agent (robot or environmental entity) in terms of observable and unobservable components:
  • Observable states: position $\mathbf{p}=[p_{x},p_{y}]$, velocity $\mathbf{v}=[v_{x},v_{y}]$, and radius $r$.
  • Unobservable states: goal position $\mathbf{g}=[g_{x},g_{y}]$, preferred speed $v_{\text{pref}}$, and heading angle $\theta$.
While acting in the environment robot knows his observable and unobservable states and also robot knows the observable states of all entities. We use the robot-centric parameterization described in Chen et al. (2017b) and Everett et al. (2018), where the robot is positioned at the origin and the x-axis is aligned with the robot’s goal direction. Denote $e_{i}$ as a type of entity $i$. The robot’s state and the $i$-th entity’s observable state at time $t$ could be represented as:
$\mathbf{s}_{t}^{r}$ $=[d_{g},v_{x},v_{y},r,v_{\text{pref}},\theta],$ (1)
$\mathbf{s}_{t}^{io}$ $=[p_{x}^{i},p_{y}^{i},v_{x}^{i},v_{y}^{i},r^{i},d^{i},r^{i}+r,e_{i}],$
where $d_{g}=\|\mathbf{p}_{t}-\mathbf{g}\|_{2}$ is the distance from the robot to the goal, and $d^{i}=\|\mathbf{p}_{t}-\mathbf{p}_{t}^{i}\|_{2}$ is the distance from the robot to the $i$-th entity.
The robot can acquire its own state and the observable states of other agents at each time step. The joint state at time $t$ is defined by:
$\mathbf{s}_{t}^{j}=[\mathbf{s}_{t}^{r},\mathbf{s}_{t}^{1o},\mathbf{s}_{t}^{2o},\ldots,\mathbf{s}_{t}^{no}].$ (2)
The robot’s velocity $\mathbf{v}_{t}$ is determined by the action command $\mathbf{a}_{t}$ from the navigation policy, i.e., $\mathbf{v}_{t}=\mathbf{a}_{t}=\pi(\mathbf{s}_{t}^{j})$.
The reward function $R(\mathbf{s}_{t}^{j},\mathbf{a}_{t})$ is defined as a mapping from the current state $\mathbf{s}_{t}^{j}$ and action $\mathbf{a}_{t}$ to a scalar value, which represents the immediate benefit of taking action $\mathbf{a}_{t}$ in state $\mathbf{s}_{t}^{j}$. Formally:
$R(\mathbf{s}_{t}^{j},\mathbf{a}_{t})=r_{t},$ (3)
where $r_{t}$ is the reward received at time step $t$.
Our objective is to find the optimal policy $\pi^{*}(\mathbf{s}_{t}^{j})$ that maximizes the expected reward.
The optimal value function is given by Sutton and Barto (2018):
$V^{*}(\mathbf{s}_{t}^{j})=\sum_{\tilde{t}=t}^{T}\gamma^{\tilde{t}\cdot v_{\text{pref}}}R_{\tilde{t}}(\mathbf{s}_{\tilde{t}}^{j},\pi^{*}(\mathbf{s}_{\tilde{t}}^{j})),$ (4)
where $\gamma\in(0,1)$ is a discount factor and $T$ is the time step at which the episode ends.
Using value iteration, the optimal policy is derived as:
$\pi^{*}\left(\mathbf{s}_{t}^{j}\right)=\arg\max_{\mathbf{a}_{t}\in\mathbf{A}}\biggl{[}R\left(\mathbf{s}_{t}^{j},\mathbf{a}_{t}\right)+\gamma^{\Delta t\cdot v_{\text{pref }}}\cdot$ (5)
$\quad\int_{\mathbf{s}_{t+\Delta t}^{j}}P\left(\mathbf{s}_{t+\Delta t}^{j}\mid\mathbf{s}_{t}^{j},\mathbf{a}_{t}\right)V^{*}\left(\mathbf{s}_{t+\Delta t}^{j}\right)d\mathbf{s}_{t+\Delta t}^{j}\biggr{]},$
where $\Delta t$ is time step duration, $R(\mathbf{s}_{t}^{j},\mathbf{a}_{t})$ is the reward function, $\mathbf{A}$ is the action space, and $P(\mathbf{s}_{t+\Delta t}^{j}\mid\mathbf{s}_{t}^{j},\mathbf{a}_{t})$ is the transition probability.
3 New Approach
In this section, we formulate our new approach which we call Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL). We propose a novel reward function. Then, we propose a model based on a neural network that utilizes the information about the types of entities. Finally, we formulate an improved version of the deep V-learning algorithm that utilizes multiprocessing and significantly improves the performance during training, validation, and testing.
3.1 Reward Function
For concise usage in the subsequent formulas we denote the entity types:
  • Adult: $A$
  • Bicycle: $B$
  • Child: $C$
  • Obstacle: $O$
Let $d(e)$ be the minimum distances between the robot and each of entity types where $e\in\{A,B,C,O\}$. Let $d_{min}=\min\{d(A),d(B),d(C),d(O)\}$ be the minimum distance between the robot and all entities around the robot over the period $[t-\Delta t,t]$. If a collision with the robot occurs for entity $e$, then $d(e)$ and $d_{min}$ are less than or equal to zero.
The function $R_{t}(t)$ represents the time reward:
$R_{t}(t)=\begin{cases}1&\text{if }t<t_{good}\\ \frac{t_{max}-t}{t_{max}-t_{good}}&\text{if }t_{good}\leq t\leq t_{max}\\ 0&\text{otherwise}\end{cases}$ (6)
where: $t_{good}$ is the time duration considered as “good" for reaching the goal, $t_{max}$ is the maximum allowable time to reach the goal.
Let $d_{max}$ be the maximum distance to the goal, which is the Euclidean distance between the starting position of the robot at time $t=0$ and the position of the goal. Let $d_{g}(t)$ be the distance to goal at time $t$. Then proximity reward can be represented as:
$R_{p}(t)=1-\frac{d_{g}(t)}{d_{max}}$ (7)
Let $R_{c}(e)$ represent the penalties for collisions. The penalties are defined as:
$R_{c}(e)=\begin{cases}-0.5,&\text{if }e=O\\ -1.0,&\text{if }e=A\\ -1.5,&\text{if }e=B\\ -2.0,&\text{if }e=C\end{cases}$ (8)
The rationale behind the collision penalties is based on the predictability of the entity’s behavior and the severity of the consequences of a collision. Static obstacles are the most predictable and have the least severe consequences, resulting in the smallest penalty (-0.5). Adults exhibit predictable behavior with moderate collision consequences, resulting in a penalty of -1.0. Cyclists, due to their higher speed and potential for severe consequences such as falling from the bicycle, incur a higher penalty (-1.5). Children are the most unpredictable and vulnerable, leading to the most severe consequences in the event of a collision, thus resulting in the highest penalty (-2.0).
Let $d_{disc}(e)$ be the discomfort distance for entity $e$:
$d_{disc}(e)=\begin{cases}0.1,&\text{if }e=A\\ 0.2,&\text{if }e=B\\ 0.2,&\text{if }e=C\end{cases}$ (9)
Let $p_{disc}(e)$ be the discomfort penalty factor for entity $e$:
$p_{disc}(e)=\begin{cases}0.5,&\text{if }e=A\\ 1.0,&\text{if }e=B\\ 1.0,&\text{if }e=C\end{cases}$ (10)
Finally, we can construct the penalty for being close to the entity:
$c_{disc}(e)=(d(e)-d_{disc}(e))*p_{disc}(e)*\Delta t$ (11)
The robot is not penalized for proximity to static obstacles in our model, since the robot does not cause discomfort for static obstacles. Therefore, the discomfort distance $d_{disc}(O)$ and discomfort penalty factor $p_{disc}(O)$ are omitted.
Let $E_{t}$ represent the entity type of the entity which is closest to the robot at time $t$. Now, our new reward function can be represented as follows:
$R(t)=\begin{cases}R_{p}(t)&\text{if }t\geq t_{max}\text{ and }\mathbf{p}_{t}\neq\mathbf{g}\\ R_{c}(E_{t})+R_{p}(t)&\text{else if }d_{min}<0\\ 1+R_{t}(t)&\text{else if }\mathbf{p}_{t}=\mathbf{g}\\ c_{disc}(\text{C})&\text{else if }0\leq d(C)<d_{disc}(C)\\ c_{disc}(\text{B})&\text{else if }0\leq d(B)<d_{disc}(B)\\ c_{disc}(\text{A})&\text{else if }0\leq d(A)<d_{disc}(A)\\ 0&\text{otherwise }\end{cases}$ (12)
The assignment of collision penalties, discomfort penalties, and discomfort distance thresholds in our model is based on the potential risk and harm associated with the robot’s interactions with different entities. For instance, a collision with a static obstacle like a fence, wall or trash can might result in minimal damage to the robot, with no risk to other entities. Conversely, collisions with children could lead to serious injury of children. Collisions with bicyclists could lead to serious injury for both the rider and the robot and could cause the rider to fall from the bicycle, increasing the potential harm. Our model, therefore, seeks to minimize these risks by training the model to avoid such collisions and dangerous situations.
3.2 Model
Our model builds upon the architecture described in Chen et al. (2019), enhancing it for the specific task of controlling a delivery robot navigating in dynamic environments with various obstacles and pedestrians. To account for different types of entities (e.g., humans, children, bicycles), we incorporated an additional module. Initially, this module uses one-hot encoding to represent each agent type. For more advanced scenarios, one could use trainable embeddings for each agent type. These encoded features or embeddings are then concatenated with the primary features of the agents and passed into the model.
We embed the state of each entity $i$ and the map tensor $M_{i}$, along with the state of the robot, into a fixed-length vector $g_{i}$ using a multi-layer perceptron (MLP):
$g_{i}=\phi_{g}(s^{r},s^{io},M_{i};W_{g})$ (13)
where $\phi_{g}(\cdot)$ is an embedding function with ReLU activations, and $W_{g}$ are the embedding weights.
The embedding vector $g_{i}$ is then fed into another MLP to obtain the pairwise interaction feature between the robot and entity $i$:
$h_{i}=\psi_{h}(g_{i};W_{h})$ (14)
where $\psi_{h}(\cdot)$ is a fully-connected layer with ReLU nonlinearity, and $W_{h}$ are the network weights.
Due to the significant variation in the number of surrounding entities across different scenes, our model must be able to handle a variable number of inputs while generating a fixed-size output. To achieve this, we employ a social attentive pooling module introduced in Chen et al. (2019), which learns the relative importance of each neighbor and the collective impact of the crowd in a data-driven manner.
The interaction embedding $g_{i}$ is transformed into an attention score $\alpha_{i}$ as follows:
$\begin{gathered}g_{m}=\frac{1}{n}\sum_{k=1}^{n}g_{k},\\ \alpha_{i}=\psi_{\alpha}(g_{i},g_{m};W_{\alpha})\end{gathered}$ (15)
where $g_{m}$ is a fixed-length embedding vector obtained by mean pooling all the individuals, $\psi_{\alpha}(\cdot)$ is an MLP with ReLU activations, and $W_{\alpha}$ are the weights.
Given the pairwise interaction vector $h_{i}$ and the corresponding attention score $\alpha_{i}$ for each neighbor $i$, the final representation of the crowd is a weighted linear combination of all pairs:
$c=\sum_{i=1}^{n}\operatorname{softmax}(\alpha_{i})h_{i},$ (16)
where $\operatorname{softmax}(\alpha_{i})=\frac{e^{\alpha_{i}}}{\sum_{j=1}^{n}e^{\alpha_{j}}}.$
Using the compact representation of the crowd $c$, we construct a planning module that estimates the state value $v$ for cooperative planning:
$v=f_{v}(s^{r},c;W_{v})$ (17)
where $f_{v}(\cdot)$ is an MLP with ReLU activations, and $W_{v}$ are the weights. We denote this value network as $V$.
3.3 Algorithm Parallel Deep V-learning
The value network is trained by the temporal-difference method with standard experience replay and fixed target network techniques. However, we improve it by using parallelism. We run several agents in several environments in parallel which helps us to collect the experience much faster than in the previous work. We call this approach “Parallel Deep V-learning", the pseudocode is shown in Algorithm 1 and Algorithm 2.
The approach that we present in Algorithms 1 and 2 is a bit similar to asynchronous methods for deep reinforcement learning Mnih et al. (2016). However, the key difference here is that we do not do asynchronous updates of neural network. Instead, we accumulate state and value pairs into Experience replay in parallel and then we make a centralized update using the whole replay buffer with random sampling. In our setup, the episode duration is much longer than the gradient learning step, so the bottleneck of the original method was in running episodes. So using parallel deep V-learning solves the problem even without asynchronous model updates. In the future, we can also test our algorithm with asynchronous model updates.
Algorithm 1 Parallel Deep V-learning
1:Run imitation learning with demonstration $D$
2:Update value network $V$ with $D$
3:Initialize target value network $\hat{V}\leftarrow V$
4:Initialize experience replay memory $E\leftarrow D$
5:Initialize parallel processes number $N$
6:while episode $<$do
7:     Initialize a multiprocessing Pool of size $N$
8:     for $k$ = episode, episode + $N$ do in parallel:
9:         Run the function $RUN\_EPISODE(k)$
10:     end for
11:     Sample random minibatch tuples from $E$
12:     Update value network $V$ by gradient descent
13:     if episode $\mod$ UpdateInterval = 0 then
14:         Update target value network $\widetilde{V}\leftarrow V$
15:     end if
16:     Update $episode\leftarrow episode+N$
17:end while
18:return $V$
Algorithm 2 RUN_EPISODE function
1:function RUN_EPISODE
2:     Initialize environment
3:     Initialize robot with current model $V$
4:     Initialize random sequence $s_{0}^{j}$
5:     Initialize temporary buffer $S$
6:     repeat
7:         $\mathbf{a}_{t}\leftarrow\begin{cases}\text{RandomAction}(),\qquad\qquad\epsilon\\ \arg\max_{\mathbf{a}_{t}\in\mathbf{A}}R(\mathbf{s}_{t}^{j},\mathbf{a}_{t})+\\ \qquad+\gamma^{\Delta t\cdot v_{\text{pref}}}V(\mathbf{s}_{t+\Delta t}^{j}),\quad 1-\epsilon\end{cases}$
8:         Make step using $\mathbf{a}_{t}$
9:         Store tuple $(s_{t}^{j},a_{t},r_{t},s_{t+\Delta t}^{j})$ in $S$
10:         Update time $t\leftarrow t+\Delta t$
11:     until $s_{t}$ is a terminal state or $t\geq t_{\max}$
12:     if Need to update memory then
13:         for $t=0$, length($S$do
14:              Get $r_{t}$, $s_{t}^{j}$ and $s_{t+1}^{j}$ from $S$
15:              Set $value_{t}=r_{t}+\gamma^{\Delta t\cdot v_{\text{pref}}}\hat{V}(s_{t+1}^{j})$
16:              Store ($s_{t}$, $value_{t}$) in $E$
17:         end for
18:     end if
19:end function
3.4 Implementation Details
We assume holonomic kinematics for the robot, i.e., it can move in any direction. The action space consists of 80 discrete actions: 5 speeds exponentially spaced between (0, $v_{pref}$] and 16 headings evenly spaced between $[0,2\pi)$.
The hidden units of functions $\phi(\cdot)$, $\psi_{h}(\cdot)$, $\psi_{\alpha}(\cdot)$, $f_{v}(\cdot)$ are (300, 200), (200, 100), (200, 200), (300, 200, 200) respectively.
We implemented the policy in PyTorch and trained it with a batch size of 100 (for each episode) using stochastic gradient descent. For imitation learning, we collected 3000 episodes of demonstration using ORCA and trained the policy 50 epochs with a learning rate of 0.01. For reinforcement learning, the learning rate is 0.001 and the discount factor $\gamma$ is 0.9. The exploration rate of the $\epsilon$-greedy policy decays linearly from 0.5 to 0.1 in the first 5000 episodes and stays at 0.1 for the remaining 5000 episodes.
4 Experiments
4.1 Simulator
Figure 1: Simulator example.
We train and test our models using our simulator. The example of one step of an episode is shown in Figure 1. The robot, adults, bicyclists, children, and static obstacles are shown. The goals of each entity (excluding static obstacles) and the robot’s goal are marked as stars. Each dynamic entity and its goal is enumerated. We use a square crossing scenario, where all the agents are randomly positioned on a square and their goal positions are chosen randomly on the opposite side of the square. Also, we use a circle crossing scenario, where all the agents are randomly positioned on a circle and their goal positions are chosen on the opposite side of the circle.
4.2 Simulation Setup
The simulated agents: adults, children, and bicyclists are controlled by Optimal Reciprocal Collision Avoidance (ORCA) van den Berg et al. (2011). The sizes and speed of agents are generated from Uniform distribution to introduce diversity. The parameters of the distribution depend on the entity type. The distribution parameters are chosen so that, on average, adults are slightly larger and faster than children, and cyclists are slightly larger and faster than adults. Additionally, we account for the presence of slow cyclists and fast children, as well as for the presence of small cyclists and large children. Consequently, it is not possible to accurately determine the type of object solely based on its size and speed.
We use a square crossing scenario, where all the agents are randomly positioned on a square of size 11m $\times$ 11m. For simplicity, in our experiments, we represent static objects as several stationary pedestrians. However, in future work, we plan to incorporate static grid processing into our method. We compare the results with the state-of-the-art method SARL Chen et al. (2019) and its modification SARL-GP which is the SARL method with a reward function that uses a Goal Proximity component.
We evaluate the effectiveness of the proposed model in the invisible setting. In this setting, the robot is invisible to other agents. This setting could be sometimes more complicated than the setting when other agents react to robot movement, as it tests the “worst case" scenario when the robot should avoid the collision itself. We also think this is a good setting for comparing methods since robots are often overlooked by humans, cyclists, and vehicles due to their size. As a result, the simulated agents react only to other agents but not to the robot.
Models are trained for 3000 episodes using imitation learning and for 50000 episodes using Parallel Deep V-learning algorithm 1. We validate models every 1024 episodes on a Validation dataset of size 500. After train step for each method, we select the model that achieved the highest reward metric on the validation dataset. We conduct a final comparison using a Test dataset consisting of 1000 episodes.
Figure 2: Collision Rate, Success Rate, Reward and Weighted Score metrics on Train and Validation data.
Figure 3: Density histograms of danger distances on Test dataset.
4.3 Results
The results of the train and validation steps are shown in Figure 2. Firstly, based on the Success Rate metric we can see that the SARL method completely fails to learn policy that can reach the goal. The SARL model can only learn a policy that successfully avoids collisions. SARL has the lowest Collision Rate metric. As our main task is to provide an algorithm that achieves the goal (while avoiding obstacles), we do not consider the SARL model in further analysis, since it fails to do such a task completely.
We can see that both methods SARL-GP and EB-CADRL (Ours) successfully converge during 50000 train episodes and get to a plateau approximately from 30000 episodes. We can see that EB-CADRL outperforms other methods both on train and validation using the Collision Rate metric and Weighted Score metric. We compute the Weighted Score metric using the following formula:
$Weighted\ Score$ $=SR-CR(A)-4.0\cdot CR(C)$ (18)
$\quad-2.0\cdot CR(B)-0.5\cdot CR(O)$
The reward metric in this case cannot be used to compare different models as each model has its own reward function. However, it indicates how the method converges and could be used to get the episode for early stopping based on results on the validation dataset.
In Table 1 we provide the result of testing each model on test data which has a size of 1000 episodes. For comparison, we use the following metrics: SR - success rate, CR - collision rate, CR(A) - collision rate with adults, CR(B) - collision rate with bicycles, CR(C) - collision rate with children, CR(O) - collision rate with static obstacles, Time - average time to reach the goal, DD(A) - mean danger distance with adults, DD(B) - mean danger distance with bicycles, DD(C) - mean danger distance with children, DD(O) - mean danger distance with static obstacles. We assume that danger starts when the distance between the robot and the entity is less than 30cm.
Method SR CR CR(A) CR(B) CR(C) CR(O) Time DD(A) DD(B) DD(C)
SARL 0 0.009 0.003 0.004 0.002 0 NaN 0.175 0.173 0.185
SARL-GP 0.681 0.207 0.027 0.062 0.029 0.089 29.32 0.13 0.113 0.131
EB-CADRL 0.681 0.068 0.02 0.018 0.002 0.028 29.45 0.118 0.154 0.165
Table 1: Metrics on Test data. We indicate best metric value in comparison between SARL-GP and EB-CADRL (since SARL fails to learn to navigate to the goal), differce less than 0.5% is not indicated.
SARL has a success rate of 0, meaning it fails to reach the goal in all attempts. Both SARL-GP and EB-CADRL have a success rate of 0.681, indicating they are equally successful in reaching the goal. SARL has the lowest collision rate, but it’s important to note that its success rate is zero, indicating it did not navigate successfully enough to encounter many obstacles. EB-CADRL has a significantly lower collision rate compared to SARL-GP, indicating better overall safety. EB-CADRL has lower collision rates with bicycles and static obstacles compared to SARL-GP and is on par with SARL in terms of collisions with children. It also has a lower collision rate with adults compared to SARL-GP, although slightly higher than SARL. SARL-GP and EB-CADRL have similar average times to reach the goal, indicating that EB-CADRL’s improved safety does not come at the expense of significantly increased travel time. The time to reach the goal for SARL is invalid since SARL has a 0 Success Rate, so it is indicated as NaN in the table. EB-CADRL maintains a safer distance from adults compared to both SARL and SARL-GP, but it is closer to bicycles and children compared to SARL-GP. However, it still achieves a balance between maintaining safe distances and minimizing collisions. To sum up, EB-CADRL is the better method overall because it achieves a good balance between reaching the goal successfully, minimizing collisions, and maintaining safe distances. It demonstrates the best safety performance while maintaining efficiency, making it suitable for robot navigation in environments with dynamic and static entities of different types.
In Figure 3 we provide the histograms of the distribution of danger distances of SARL-GP and EB-CADRL methods. We can see that the EB-CADRL model tends to be safer with bicyclists and children.
4.4 Ablation Experiments
To evaluate the impact of including the agent type as a feature, we conducted an ablation study by training and testing two models: one incorporating the agent type and the other excluding it. While the model can infer some aspects of the agent type through features such as radius and velocity, these features alone do not provide a clear distinction between different agent types. For instance, a small, slow-moving cyclist or a large, fast-moving human could mislead the model. Therefore, we hypothesize that explicitly adding the agent type to the feature set could enhance the model’s performance. Both models in this experiment utilize a new reward function and are based on the EB-CADRL method.
For ablation experiments, we use circle and square crossing scenarios. The radius of the circle is 6m and the size of the square is 13m $\times$ 13m.
To test models during ablation experiments we use the following metrics. Reward is the reward function that we formulated in 3.1. Success measures the percentage of times the robot successfully reached its goal without collisions. Collision metrics (Adult, Bicycle, Child, Obstacle) measure the frequency of collisions with different entities. Danger time is the mean time during which the robot was in a potentially dangerous state (close to collision). A lower value means the robot spent less time in dangerous zones in which there is a risk of getting into a collision. Apart from that, the long presence of the robot in such areas may cause discomfort to surrounding pedestrians and other entities which is not desirable. Time is the average time it took for the robot to reach its goal. A lower time indicates that the robot was able to complete its task more quickly.
No entity type With entity type
Reward 0.037 0.052
Success 0.448 0.501
Collision (Bicycle) 0.004 0.002
Collision (Child) 0.001 0.001
Collision (Human) 0.015 0.016
Collision (Obstacle) 0.002 0.005
Danger time 0.913 0.746
Time 39.078 32.857
Table 2: Metrics on Test data during ablation experiments.
The results of the ablation experiment on the Test dataset are shown in Table 2. The reinforcement learning model with entity type outperforms the one without entity type in several key aspects. The model with entity type achieved a higher reward and a better success rate, indicating superior overall performance. It also registered fewer collisions with bicycles, a shorter time to reach the goal, and less time in danger. However, it is worth noting that the model without entity type had fewer collisions with adults and obstacles. Overall, the inclusion of entity type in the model appears to enhance its performance in most aspects.
Figure 4: Reward and success rate metrics on Train and Validation data during ablation experiments.
5 Discussion
While our EB-CADRL method shows promising results, several limitations remain. The simulation environment may not fully capture real-world complexities such as varying terrains and unpredictable human behaviors. The predefined entity types and simplified reward function might limit the model’s adaptability to more nuanced scenarios. Moreover, the current implementation assumes holonomic kinematics, which may not be applicable to all types of robots. Future work should focus on real-world testing, incorporating advanced perception systems, developing adaptive reward functions, optimizing computational efficiency, extending the model to non-holonomic robots, and exploring multi-robot coordination. Addressing these areas will enhance the robustness, safety, and efficiency of autonomous robot navigation in dynamic environments.
6 Conclusion
In this study, we introduced a novel reinforcement learning approach, termed Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL), for the navigation of autonomous delivery robots in dynamic environments. Our primary focus was on enhancing safety and efficiency by leveraging entity-specific information to guide the robot’s interactions with different types of agents and obstacles. Our proposed reward function penalized the robot differently based on the type of entity involved in a collision and its proximity to entities, thereby ensuring more context-aware and safer navigation. Our experimental results demonstrate that EB-CADRL significantly outperforms traditional navigation methods and state-of-the-art techniques. Notably, our approach achieved a higher success rate in reaching the goal while maintaining a lower collision rate across various entity types. The model effectively balanced the trade-off between safety and efficiency, maintaining safe distances from entities without significantly increasing travel time. The ablation experiments further validated the importance of incorporating entity-specific information, as models with this feature consistently performed better in terms of reward, success rate, and overall safety metrics compared to those without it. This underscores the relevance and added value of our work in developing more sophisticated and context-aware navigation systems for autonomous robots. In conclusion, our proposed EB-CADRL approach represents a significant advancement in the field of autonomous robot navigation, offering a robust, safe, and efficient solution for navigating dynamic, crowded environments.
Appendix A Appendix
In this section we show figures of collision rate for each entity type on train and validation data.
Figure 5: Collision Rate by Entity Type on Train and Validation data.